OpenStudy (anonymous):

How do you graph this straight line on a graph y<_(-4/5)x+4

OpenStudy (anonymous):

\[y \le \frac{-4}{5}x+4\] Is the region of all the points under and on the line \[y = \frac{-4}{5}x+4\]

OpenStudy (anonymous):

below the line